Let R br a realtion defined from the set A={1,2,3,4} to the set B={2,3,4,5}as a€A, b€B aRb <-> a+b=5




1. What are the orderd pairs in the relation R




2. Represent R with a matrix

1
Expert's answer
2022-01-19T17:58:39-0500

1) Since "1+4=5" then "\left( {1,4} \right) \in R" .Similarly we get "(2,3) \in R,\,(3,2) \in R" .

Then "R = \{ \left( {1,4} \right),\,(2,3),(3,2)\}" .

Answer: "R = \{ \left( {1,4} \right),\,(2,3),(3,2)\}"

2) Let's build a matrix of relation.

If the elements are in a relationship, then write 1, otherwise - 0. We have: